Pharmax Distributors specializes in the distribution of pharmaceutical products at the retail level. Pharmax's customer base is comprised of three main retail groups: general supermarket chains, large drugstore chains, and small, single-store "Ma & Pa" pharmacies. To evaluate the relative profitability of each customer group, Pharmax plans to use an allocation system that recognizes the relative demands that these groups place on the company's resources. Parmax's chief accountant examined the firm's support costs and activities and collected the following data for a typical operating period:

Support Activity

Budgeted Cost

Total Cost Driver
Order processing $86,000 2,110 customer orders
Line item processing $63,000 24,460 line items
Store deliveries $75,000 1,520 deliveries
Carton handling $74,000 80,400 cartons shipped


She also gathered the following data for a typical operating period:

General Supermarket

Drugstore Chains Ma & Pa
Customer orders 140 370 1,600
Line items 1,680 5,180 17,600
Deliveries 120 300 1,100
Cartons shipped 33,600 27,000 19,800



REQUIRED [ROUND ALL ALLOCATION RATES TO THE NEAREST CENT AND ALLOCATIONS TO THE NEAREST DOLLAR.]

Part A (5 tries; 5 points)
If Pharmax uses a single-driver system to allocate its support costs, with number of deliveries as the single driver, how much would be allocated to its Drugstore Chain customers? $58,815


Part B (5 tries; 5 points)
If Pharmax uses an activity-based costing system to allocate its overhead costs, how much would be allocated to its General Supermarket customers?

137,249 is not correct
Answer #1

Calculate Following

Activity Activity rate Driver Allocated cost to General supermarket customers
Order processing 86000/2110 = 40.76 140 5706
Line item processing 63000/24460 = 2.58 1680 4334
Store deliveries 75000/1520 = 49.34 120 5921
Carton handling 74000/80400 = 0.92 33600 30921
Total 46873